OpenStudy (anonymous):

find the derivative of the hyperbolic of 1/4sinh(2x)-x/2, i know i have to use the product rules but i am confused on what to use with the product rule

OpenStudy (anonymous):

not sure if you mean( 1/4(sinh 2x)) - (x/2) If yes it's( (1/2) cosh (2x)) - (1/2)

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\sinh(2x)-\frac{ x }{ 2 }\] this is what it is, and it is a hyperbolic derivative

OpenStudy (anonymous):

Yes then that is the answer. Disecting the problem: \[\frac{ 1 }{ 4} \sin h ( 2x) = \frac{ 1 }{ 4 } ( 2) \cos h (2x)\] and d/dx of - x/2 = 1/2 so the answer : \[\frac{ 1 }{ 2 } \cos h (2x) - \frac{ 1 }{ }\]

OpenStudy (anonymous):

-1/2 *

OpenStudy (anonymous):

uhhh, the book says sinh^2(x)

OpenStudy (anonymous):

Are you sure? I even checked with an online calc to make sure all my work is right. By no means should you use this tool to solve all probs lol http://www.derivative-calculator.net/#expr=1%2F4%20sinh%20%282x%29%20-%20x%2F2

OpenStudy (anonymous):

click go

OpenStudy (anonymous):

it might be major simplifcation your answer is probabaly right